OpenStudy (anonymous):

Look at the following number sequence. 3,7,11,15 based on the pattern what are the two next numbers?

OpenStudy (texaschic101):

7 - 3 = 4 11 - 7 = 4 15 - 11 = 4 you see the pattern....the difference is 4 so what is 15 + 4 ? and what is the answer of that, plus 4 ?